OpenStudy (anonymous):

MEDAL AND FAN John puts the letters of the word SCIENCE on pieces of paper and puts the pieces in a bag. He draws two letters without looking. The first piece of paper is not replaced before the second piece is drawn. What is the probability that John draws a C twice?

OpenStudy (anonymous):

2 out of 7. if you're expecting a more mathmatical answer, i can't give you that. :P

OpenStudy (anonymous):

thats not an answer choice tho.. A) 2/49 B)1/7 C)1/14 D)1/21

OpenStudy (anonymous):

you should of given the possibilties of answers. sorry. :(

OpenStudy (anonymous):

do those make sense now?

OpenStudy (anonymous):

uhhh im not sure...

OpenStudy (anonymous):

do you know anyone else that can helpp?

OpenStudy (kirbykirby):

There are 2 "C" out of 7 letters. P("C" on first try AND "C" on 2nd try) = P("C" on first try") * P("C" on 2nd try) P("C" on first try) = 2/7 On the 2nd try, you removed a C, so only 6 letters are left. Hence, P("C" on 2nd try) = 1/6 Thus,P("C" on first try AND "C" on 2nd try) = (2/7) * (1/6) = 2/42 = 1/21